OpenStudy (anonymous):

Jennifer is a safety inspector for a family gym. Recently, guests have been complaining about the diving board saying that it is leaning towards the water. Jennifer needs to know whether this diving board is standing up straight or not. Jennifer measures the height of the diving board to be 9 feet. The pool is 12 feet across and she measures the diagonal from the other side of the pool to the top of the diving board to be 15 feet. Is the diving board at a 90 degree angle with the ground? Explain how you know.

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

is there a picture that go along with this problem? i can't seem to make a picture of this situation... can you sketch it for me?

OpenStudy (anonymous):

@dpaInc

OpenStudy (anonymous):

yes it is... this is a utilization of the converse of the pythagorean theorem in that if the sides of a triangle satisfy a^2 + b^2 = c^2, then it is a right triangle.... since 9^2 + 12^2 = 15^2, the board is perpendicular to the ground...